# DisastIR: A Comprehensive Information Retrieval Benchmark for Disaster Management

### v1

@@ Appendix I Additional Analyses of Model Performance across 48 Tasks
Biological 64.38 62.94 43.89 57.80 40.49 69.47 56.50
Chemical 64.29 60.87 42.16 59.77 39.28 70.09 56.08
Environmental 65.73 60.71 42.67 59.05 43.81 69.91 56.98
Extraterrestrial 65.50 65.11 38.38 60.94 38.15 69.16 56.21
Geohazard 62.01 60.48 40.11 56.35 40.05 67.06 54.34
Meteorological&hydrological 64.65 62.91 39.82 57.32 38.05 67.04 54.97
Societal 65.72 65.82 39.30 58.65 41.65 68.60 56.62
Technological 64.26 64.83 38.31 57.85 40.30 69.16 55.78
Avg. 64.57 62.96 40.58 58.47 40.22 68.81 55.93
@@
Biological 69.88 63.74 57.69 62.45 47.40 68.75 61.65
Chemical 69.31 56.86 57.02 59.12 45.99 66.66 59.16
Environmental 69.14 63.86 57.85 61.98 52.46 68.36 62.28
Extraterrestrial 65.64 62.43 53.94 57.93 45.98 69.16 59.18
Geohazard 66.26 62.29 55.90 58.56 48.33 65.33 59.44
Meteorological&hydrological 69.42 63.50 58.28 61.66 47.69 68.60 61.52
Societal 68.68 66.72 57.45 58.88 50.42 67.42 61.59
Technological 68.59 64.78 56.57 59.77 48.24 68.62 61.10
Avg. 68.36 63.02 56.84 60.04 48.31 67.86 60.74
@@
Biological 76.42 69.25 42.42 69.75 58.26 77.86 65.66
Chemical 74.82 69.52 42.07 68.68 57.96 76.22 64.88
Environmental 75.31 68.66 43.82 69.63 61.09 77.21 65.95
Extraterrestrial 73.06 70.08 44.88 68.30 56.31 78.18 65.14
Geohazard 73.91 67.35 41.35 66.72 57.88 75.91 63.85
Meteorological&hydrological 75.33 69.91 43.11 68.37 58.88 76.85 65.41
Societal 75.38 72.11 44.57 68.78 61.06 76.54 66.41
Technological 73.91 71.01 43.18 68.85 58.42 77.31 65.45
Avg. 74.77 69.74 43.18 68.64 58.73 77.01 65.34
sentence-transformers-all-mpnet-base-v2
Biological 14.00 7.64 17.18 43.65 27.58 35.38 24.24
Chemical 13.09 9.93 14.53 48.81 24.14 35.84 24.39
Environmental 12.45 10.01 18.17 47.32 30.32 35.48 25.62
Extraterrestrial 16.24 8.11 14.82 48.79 27.48 42.02 26.24
Geohazard 14.96 7.75 13.30 43.82 27.70 36.33 23.98
Meteorological&hydrological 16.52 7.84 16.39 44.34 28.07 36.99 25.02
Societal 20.10 14.49 18.41 45.98 27.31 39.06 27.56
Technological 13.16 12.41 16.55 47.68 26.44 36.75 25.50
Avg. 15.06 9.77 16.17 46.30 27.38 37.23 25.32
@@
Biological 10.09 5.90 45.04 61.49 44.70 49.42 36.11
Chemical 9.73 5.62 40.67 62.27 41.06 53.75 35.52
Environmental 6.03 4.75 38.62 63.00 48.95 40.21 33.59
Extraterrestrial 10.42 4.54 33.58 59.45 41.39 44.64 32.34
Geohazard 7.96 3.33 34.20 58.54 41.68 43.00 31.45
Meteorological&hydrological 9.16 5.31 38.46 60.76 40.55 43.60 32.97
Societal 10.00 7.88 39.76 60.39 44.40 48.17 35.10
Technological 10.39 5.44 38.00 60.50 40.08 50.33 34.13
Avg. 9.22 5.35 38.54 60.80 42.85 46.64 33.90

### v2

@@ Appendix I Performance of Evaluated Models
Biological 64.21 62.79 43.11 57.82 40.36 68.66 56.16
Chemical 64.22 60.92 41.70 59.77 39.37 69.16 55.86
Environmental 65.66 60.40 41.92 58.91 43.62 69.31 56.64
Extraterrestrial 65.29 64.74 37.89 60.75 38.84 68.20 55.95
Geohazard 61.90 60.32 39.84 55.92 40.44 66.14 54.09
Meteorological&hydrological 64.24 62.74 39.44 57.19 37.94 66.12 54.61
Societal 65.49 65.63 38.81 58.36 41.72 67.70 56.29
Technological 63.97 64.75 37.88 57.70 39.80 68.42 55.42
Avg. 64.37 62.79 40.07 58.30 40.26 67.96 55.63
@@
Biological 69.72 63.54 56.65 62.33 47.93 68.08 61.37
Chemical 69.10 56.97 56.36 59.00 45.84 65.88 58.86
Environmental 69.01 63.61 57.10 61.78 52.05 67.91 61.91
Extraterrestrial 65.45 62.21 53.16 57.79 45.98 68.42 58.84
Geohazard 66.10 62.08 55.43 58.16 48.53 64.47 59.13
Meteorological&hydrological 69.13 63.37 57.63 61.52 47.61 67.50 61.13
Societal 68.40 66.54 56.95 58.62 50.21 66.67 61.23
Technological 68.21 64.63 55.75 59.57 47.84 67.94 60.66
Avg. 68.14 62.87 56.13 59.85 48.25 67.11 60.39
@@
Biological 76.22 69.00 41.41 69.62 58.10 77.03 65.23
Chemical 74.61 69.41 41.58 68.47 57.64 75.33 64.51
Environmental 75.20 68.39 43.19 69.37 60.63 76.54 65.55
Extraterrestrial 72.83 69.73 44.26 68.16 56.37 77.37 64.79
Geohazard 73.78 67.16 40.89 66.24 58.05 74.93 63.51
Meteorological&hydrological 75.03 69.71 42.59 68.16 58.49 75.67 64.94
Societal 75.06 71.88 44.08 68.51 60.55 75.63 65.95
Technological 73.68 70.83 42.43 68.62 57.26 76.50 64.89
Avg. 74.55 69.51 42.55 68.39 58.39 76.13 64.92
thenlper-gte-small
Biological 18.78 12.74 13.26 51.17 40.04 55.52 31.92
Chemical 18.64 12.89 12.90 53.93 35.96 54.87 31.53
Environmental 13.65 7.32 8.02 47.89 42.45 53.68 28.83
Extraterrestrial 19.50 6.40 10.64 50.69 36.66 57.62 30.25
Geohazard 16.16 5.30 9.30 45.85 36.16 53.58 27.72
Meteorological&hydrological 17.88 8.92 11.34 49.99 37.82 55.61 30.26
Societal 20.18 9.82 11.12 46.59 36.93 55.15 29.96
Technological 20.84 9.26 11.74 52.41 35.69 57.71 31.27
Avg. 18.20 9.08 11.04 49.81 37.71 55.47 30.22
@@
Biological 9.95 5.91 44.29 61.38 44.74 48.79 35.84
Chemical 9.69 5.57 40.17 62.04 41.18 53.04 35.28
Environmental 6.01 4.73 38.02 62.79 48.67 39.85 33.35
Extraterrestrial 10.36 4.50 33.16 59.23 41.91 44.01 32.19
Geohazard 7.93 3.33 33.93 58.13 41.61 42.19 31.19
Meteorological&hydrological 9.15 5.33 38.10 60.63 40.33 42.71 32.71
Societal 9.95 7.88 39.41 60.21 44.04 47.51 34.83
Technological 10.22 5.44 37.40 60.22 39.95 49.82 33.84
Avg. 9.16 5.34 38.06 60.58 42.80 45.99 33.65